Brighton Services repairs locomotive engines. It employs 100 full-time workers at $14 per hour. Despite operating at capacity, last year's performance was a great disappointment to the managers. In total, 10 jobs were accepted and completed, incurring the following total costs.

Direct materials $ 1,045,400
Direct labor 4,200,000
Manufacturing overhead 975,000

Of the $975,000 manufacturing overhead, 40 percent was variable overhead and 60 percent was fixed.

This year, Brighton Services expects to operate at the same activity level as last year, and overhead costs and the wage rate are not expected to change. For the first quarter of this year, Brighton Services completed two jobs and was beginning the third (Job 103). The costs incurred follow.

Job Direct Materials Direct Labor
101 $ 138,200 $ 490,000
102 103,000 313,000
103 95,000 195,200
Total manufacturing overhead 272,200
Total marketing and administrative costs 120,000

You are a consultant associated with Lodi Consultants, which Brighton Services has asked for help. Lodi's senior partner has examined Brighton Services's accounts and has decided to divide actual factory overhead by job into fixed and variable portions as follows.

Actual Manufacturing Overhead
Variable Fixed
101 $ 30,900 $ 105,000
102 28,500 89,200
103 5,600 13,000
$ 65,000 $ 207,200

In the first quarter of this year, 40 percent of marketing and administrative cost was variable and 60 percent was fixed. You are told that Jobs 101 and 102 were sold for $855,000 and $570,000, respectively. All over- or underapplied overhead for the quarter is written off to Cost of Goods Sold.

Required:

a. Present in T-accounts the actual manufacturing cost flows for the three jobs in the first quarter of this year.

b. Using last year's overhead costs and direct labor-hours as this year's estimate, calculate predetermined overhead rates per direct labor-hour for variable and fixed overhead.

c. Present in T-accounts the normal manufacturing cost flows for the three jobs in the first quarter of this year. Use the overhead rates derived in requirement (b).

d. Calculate operating profit (loss) for the first quarter of this year under actual and normal costing systems.

Present in T-accounts the actual manufacturing cost flows for the three jobs in the first quarter of this year. Materials InvPredetermined Overhead Rate (Per Direct Labor-Hour):

Variable overhead rate$1.30

Fixed overhead rate$1.95

Calculate operating profit (loss) for the first quarter of this year under actual and normal costing systems. (Round your final answers to nearest whole dollar amounts. Loss amounts should be indicated with a minus sign.)

Materials Inventory Wages Payable
Beg. Bal. 0 0 Beg. Bal. 0 0
0 138,200 101: DM 0 490,000 101: DL
0 103,000 102: DM 0 313,000 102: DL
0 95,000 103: DM 0 195,200 103: DL
0 0 0 0
End. Bal. 336,200 End. Bal. 998,200
Variable Manufacturing Overhead Fixed Manufacturing Overhead
Actual 65,000 ??? 101: Variable Actual ??? ??? 101: Fixed
Overapplied ??? ??? 102: Variable 0 ??? 102: Fixed
0 ??? 103: Variable 0 ??? 103: Fixed
0 0 0 0
End. Bal. 65000 0 ??? Underapplied
End. Bal.
Work-in-Process Inventory Finished Goods Inventory
Beg. Bal. 0 0 Beg. Bal. 0 0
??? ??? ??? Total 101 Finished Goods Total 101 Finished Goods ??? ??? Cost of Goods Sold
??? ??? ??? Total 102 Finished Goods Total 102 Finished Goods ??? 0
??? ??? 0 0 0
??? ??? 0 End. Bal.
0 0
End. Bal.
Cost of Goods Sold Under-or Overapplied Overhead
Beg. Bal. ??? 0 Beg. Bal. 0 0
Finished Goods ??? 0 ??? ??? ??? ???
0 0 0 0
End. Bal. End. Bal.
Actual Normal
Operating profit (loss) ??? ???
Answer #1

Answer-c:

Materials Inventory Wages Payable
Beg. Bal. 0 0 Beg. Bal. 0 0
0 1,38,200 101: DM 0 4,90,000 101: DL
0 1,03,000 102: DM 0 3,13,000 102: DL
0 95,000 103: DM 0 1,95,200 103: DL
0 0 0 0
End. Bal. 3,36,200 End. Bal. 9,98,200
Variable Manufacturing Overhead Fixed Manufacturing Overhead
Actual 65,000 45,500 101: Variable Actual 2,07,200 68,250 101: Fixed
Overapplied 27,690 29,064 102: Variable 0 43,596 102: Fixed
0 18,126 103: Variable 0 27,189 103: Fixed
0 0 0            68,165 Underapplied
End. Bal. 0 0
End. Bal.
Work-in-Process Inventory Finished Goods Inventory
Beg. Bal. 0 0 Beg. Bal. 0 0
Total DM 3,36,200 3,00,950 Total 101 Finished Goods Total 101 Finished Goods 3,00,950 7,89,610 Cost of Goods Sold
Total DL 9,98,200 4,88,660 Total 102 Finished Goods Total 102 Finished Goods 4,88,660 0
Total Fixed MOH 92,690 0 0 0
Total Variable MOH 1,39,035 0 End. Bal.
0 0
End. Bal. 7,76,515
Cost of Goods Sold Under-or Overapplied Overhead
Beg. Bal. 0 0 Beg. Bal. 0 0
Finished Goods 7,89,610 0 Under-applied 68,165 27,690 Over-applied
0 0 0 0
End. Bal. 7,89,610 End. Bal. 40,475

Total direct labor hor = $4,200,000/$14 = 300,000 DLH

Total Variable Manufacturing Overhead Applied:
Job DLH Variable Mfg
OH per DLH
Variable
Mfg OH
applied
101 $490,000/14 = 35,000 $       1.30 $       45,500
102 $313,000/14 = 22,357 $       1.30 $       29,064
103 $195,200/14 = 13,943 $       1.30 $       18,126
Total $       92,690
Total Fixed Manufacturing Overhead Applied:
Job DLH Fixed Mfg
OH per DLH
Variable
Mfg OH
applied
101 $490,000/14 = 35,000 $       1.95 $    68,250
102 $313,000/14 = 22,357 $       1.95 $    43,596
103 $195,200/14 = 13,943 $       1.95 $    27,189
Total $1,39,035

Total 101 Finished Goods = $138,200 + 490,000 + 45,500 + 68,250 = $300,950

Total 102 Finished Goods = $103,000 + 313,000 + 29,064 + 43,596 = $488,660

Answer-d:

Actual Normal
Sales revenue $        1,425,000 $        1,425,000
Less: Cost of goods sold            1,297,800                789,610
Less: Under-applied overhead                  40,475
Gross profit               127,200                594,915
Marketing and administrative expense               120,000                120,000
Operating profit $               7,200 $           474,915